OpenStudy (anonymous):

Which pair of ratios is proportional? a 5/7 and 25/35 b 6/9 and 18/36 c 9/18 and 16/36 d 35/120 and 75/240

OpenStudy (igreen):

5 / 7 = ? 25 / 35 = ?

OpenStudy (anonymous):

b

OpenStudy (igreen):

@kelsey1825 Don't give answers, plus your answer is wrong. Please read the code-of-conduct: http://openstudy.com/code-of-conduct

OpenStudy (midhun.madhu1987):

Equal ratios in the simplest form are called Proportion.

OpenStudy (anonymous):

yes:

OpenStudy (midhun.madhu1987):



OpenStudy (igreen):

Can you divide both of these? What do you get? 5 / 7 = ? 25 / 35 = ?

OpenStudy (anonymous):

they are both 1.4

OpenStudy (igreen):

Well, you divided them wrong but it doesn't matter.. So yes, they are proportional!

OpenStudy (anonymous):

lol sorry

OpenStudy (midhun.madhu1987):

Take an example... option c: